OpenStudy (anonymous):

simplify each expression sqrt72n^7

OpenStudy (saifoo.khan):

Try this one!

OpenStudy (anonymous):

i did but im stuck -___-

OpenStudy (saifoo.khan):

Where?

OpenStudy (anonymous):

n^7

OpenStudy (anonymous):

how do i simplify it

OpenStudy (saifoo.khan):

\[\Large n^\frac72 \to n^\frac62 \times n^\frac12 \to n^3 \times n^\frac12\]Therefore, \[n^3 \sqrt n\]Simple enough? :D

OpenStudy (anonymous):

yes so the answer is 3n^3 sqrtn

OpenStudy (saifoo.khan):

Now, you didn't solved sqrt(72) correctly.

OpenStudy (anonymous):

i dont got a calculator with me